1. A project has an initial cost of $59,925, expected net cash inflows of $14,000 per...

1.

A project has an initial cost of $59,925, expected net cash inflows of $14,000 per year for 6 years, and a cost of capital of 9%. What is the project's PI? Do not round your intermediate calculations. Round your answer to two decimal places.

2.

A project has an initial cost of $56,300, expected net cash inflows of $12,000 per year for 8 years, and a cost of capital of 12%. What is the project's payback period? Round your answer to two decimal places.

How many years ____

3.

A project has an initial cost of $60,000, expected net cash inflows of $14,000 per year for 7 years, and a cost of capital of 13%. What is the project's discounted payback period? Round your answer to two decimal places.

How many years _____
Answer #1

1.PI = Present value of cash inflows/Initial Cost

= 14,000*PVAF(9%, 6 years)/59,925

= 14,000*4.4859/59,925

= 1.05

2.Payback period = Initial cost/Annual cash inflows

= 56,300/12,000

= 4.69 years

3.Discounted payback period:

Year

Cash flow

PVF

Discounted cash flow

Cumulative Cash flow

0

-60,000

1

-60,000

-60,000

1

14,000

0.885

12,390

-47,610

2

14,000

0.783

10,962

-36,648

3

14,000

0.693

9,702

-26,946

4

14,000

0.613

8,582

-18,364

5

14,000

0.543

7,602

-10,762

6

14,000

0.480

6,720

-4,042

7

14,000

0.425

5,950

1,908

Discounted payback period = 6 + 4,042/5,950

= 6.68 years